The amount of time (measured in days) a watch will run without having to be reset is an exponential random variable with λ = 1/50. Find the probability that

  1. A watch will have to be reset in less than 20 days
  2. A watch will go more than 60 days before a reset
Answer #1

P(X < x) = 1 - e-x/50

a) P(X < 20) = 1 - e-20/50 = 0.3297

b) P(X > 60) = e-60/50 = 0.3012